# Joint Value Estimation and Bidding in Repeated First-Price Auctions

### v1

@@ 4.2 Regret Upper Bound
The following theorem establishes the regret upper bound of Algorithm 3.
Theorem 4.3 (Upper bound of Algorithm 3).
Let $\pi$ be the resulting bidding strategy of Algorithm 3 (combined with Algorithm 2). Then under Assumptions 2.12.3 and 2.5, the policy $\pi$ achieves an expected regret
@@
It is clear that Theorem 4.3 implies Theorem 2.6. The remainder of this section is devoted to the proof of Theorem 4.3. Since all events with probability $O(T^{-1})$ only contribute $O(S)$ regret in total, WLOG we assume that the guarantees in Assumption 2.5 and Lemma 4.1 hold almost surely. Adopting the notation [^1] Warning: not to confuse $\widehat{r}_{t}(b)$ with $\widehat{r}_{t}^{(s)}(b)$ in Algorithm 3: $\widehat{r}_{t}(b)$ is the reward w.r.t. $\widehat{G}_{t}(b)$, while $\widehat{r}_{t}^{(s)}(b)$ further plugs in the estimated parameters from stage $s$ for $\theta_{*,0}$ and $\theta_{*,1}$.
@@
In Algorithm 3, for every $t\in[T]$ and $s\in[S]$, it holds that $\max_{b^{*}\in{\mathcal{B}}}\widehat{r}_{t}(b^{*})-\widehat{r}_{t}(b)\leq 8\cdot 2^{-s}$ for all $b\in B_{s}$.
In other words, Lemma 4.4 states that the any bid that has survived into stage $s$ is $O(2^{-s})$-close to optimal under $\widehat{r}_{t}$. Consequently, if we denote by $\Phi^{(s)}\subseteq[T]$ the set of time indices belonging to stage $s\in[S]$, and $\Phi^{(S+1)}\subseteq[T]$ the set of time indices where all widths are below $\frac{1}{\sqrt{T}}$, we get
@@
Here (a) uses $|{\mathbb{E}}[r_{t}(b)]-\widehat{r}_{t}(b)|\leq\|G_{t}-\widehat{G}_{t}\|_{\infty}\leq\varepsilon_{t}$ by (2), (10), and Assumption 2.5, (b) follows from Lemma 4.4, and (c) uses $S=\lceil\log_{2}\sqrt{T}\rceil$ and $w_{t}^{(s)}(b_{t})\geq 2^{-s}$ whenever $t\in\Phi^{(s)}$. Then Theorem 4.3 is a direct consequence of the following upper bound, which is essentially the elliptical potential lemma applied to the sum $\sum_{t\in\Phi^{(s)}}w_{t}^{(s)}(b_{t})$.
@@ 5.1 An Adaptive Bid Elimination Algorithm
where $\delta_{t}$ is the concentration parameter in Assumption 2.7 which is assumed to be known. Unlike $\widehat{e}_{t}(b)$, the truncated estimator $\widetilde{e}_{t}(b)$ is biased. The following lemma summarizes the bias-variance properties of $\widetilde{e}_{t}(b)$.
@@
In the sequel, we will use $\sigma_{t}(b)^{2}$ in (23) as the variance proxy for the truncated IPW estimator $\widetilde{e}_{t}(b)$. In particular, since $\widehat{G}_{t}(b)$ could be close to 00 or $1$ (i.e. no overlap condition), the quantity $\sigma_{t}(b)$ can indeed become prohibitively large.
@@ 5.1.2 Base algorithm
1 Input: Time indices $\Phi_{t}\subseteq[t-1]$, bid space $B_{t}\subseteq{\mathcal{B}}$, estimated CDF $\widehat{G}_{t}$, time horizon $T$.
@@
7Set $\gamma\leftarrow\lambda+c_{1}\log(2T)+c_{2}\sqrt{\sum_{\tau\in\Phi_{t}}\delta_{\tau}^{2}}$, with $\delta_{\tau}$ given in Assumption 2.7 and absolute constants $c_{1},c_{2}>0$ defined in Lemma A.6.
@@
9       Compute width $w_{t}(b)\leftarrow c_{3}L\gamma\left(\widehat{G}_{t}(b)(1-\widehat{G}_{t}(b))\|x_{t}\|_{A_{t}^{-1}}+\gamma\|x_{t}\|_{A_{t}^{-1}}^{2}\right)$ with absolute constant $c_{3}>0$ defined in Lemma A.7.
@@
Next we present the base algorithm LIN-TE in Algorithm 4 based on $\widetilde{e}_{t}(b)$. Again, we will apply linear regression of $\widetilde{e}_{t}(b_{t})$ against $x_{t}$ to estimate $\theta_{*}$, but Lemma 5.1 shows that the ordinary least squares will suffer from heteroskedasticity. To address this problem, we use the following weighted least squares estimator:
@@
where $\gamma$ is defined in Line 7 of Algorithm 4.
As for the quantities $u_{t,0}(b)$ and $u_{t,1}(b)$ in Lines 10 and 11, note that the high-probability event in Lemma 5.2 yields that for every $b\in{\mathcal{B}}$,
@@ 5.1.3 Master algorithm
1 Input: Time horizon $T$, bid space ${\mathcal{B}}=[0,1]$, estimation oracle ${\mathcal{O}}$.
2Initialize: set $S=\lceil\log_{2}\sqrt{T}\rceil$, $\Phi^{(s)}_{1}=\varnothing$ for $s\in[S]$.
@@
Similar to Algorithm 3, the master algorithm in Algorithm 5 partitions $[t]$ into multiple stages to ensure the conditional independence condition in Lemma 5.2. However, the main distinct feature of Algorithm 5 lies in the selection from two different criteria $u_{t,0}$ and $u_{t,1}$. The reason behind this selection is that, the rewards
@@
in (14) and (15) are sensitive to the estimation error of $\theta_{*}^{\top}x_{t}$ in different ways: in $\widehat{r}_{t,0}(b)$ the leading coefficient is $\widehat{G}_{t}(b)$, while in $\widehat{r}_{t,1}(b)$ the leading coefficient has magnitude $1-\widehat{G}_{t}(b)$. Therefore, a possibly large estimation error of $\theta_{*}^{\top}x_{t}$ when $\widehat{G}_{t}(b)$ is close to 00 (resp. $1$) does not transform into a large estimation error in $\widehat{r}_{t,0}(b)$ (resp. $\widehat{r}_{t,1}(b)$); this is essentially a “variance reduction” step. The key then is to choose the better reward ($\widehat{r}_{t,0}(b)$ or $\widehat{r}_{t,1}(b)$) for decision making, depending on whether $\widehat{G}_{t}(b)$ is close to 00 or $1$ for approximately optimal bids $b$.
To this end, we compute the maximizer $b_{*,1}$ of the map $b\mapsto u_{t,1}(b)$, and check if $\widehat{G}_{t}(b_{*,1})$ is close to 00 or $1$. A small $\widehat{G}_{t}(b_{*,1})$ suggests a small CDF and thus the choice of $\widehat{r}_{t,0}$, and a large $\widehat{G}_{t}(b_{*,1})$ suggests a big CDF and the other option $\widehat{r}_{t,1}$. This is the criterion $i\in\{0,1\}$ in Line 9 of Algorithm 5. The following result shows that the selected reward $\widehat{r}_{t,i}(b)$ enjoys a small width $w_{t}(b)$ that scales with $\widehat{G}_{t}(b)(1-\widehat{G}_{t}(b))=\sigma_{t}(b)^{-2}$ for all candidate bids.
@@
Suppose the event in Lemma 5.2 holds. Let $\widehat{r}_{t,0}$ and $\widehat{r}_{t,1}$ be defined as in (14) and (15). For any $s\in[S]$ and for the criterion $i\in\{0,1\}$ selected in Algorithm 5, it holds that
@@
where $b^{(s)}_{*,i}=\operatorname*{arg\,max}_{b\in B_{s}}u^{(s)}_{t,i}(b)$ for $i=0,1$ as defined in Algorithm 4.
@@ 5.2 Regret Upper Bound
The regret upper bound of Algorithm 5 is established in the following result, which implies Theorem 2.8.
Theorem 5.4 (Upper bound of Algorithm 5).
Let $\pi$ be the resulting bidding strategy of Algorithm 5 (combined with Algorithm 4). Then under Assumptions 2.12.3 and 2.7, the policy $\pi$ achieves an expected regret
@@
Again, we present the proof of Theorem 5.4 in the remainder of this section. WLOG we assume Assmuption 2.7 and the high-probability event in Lemma 5.2 holds almost surely. We also assume that $\delta_{t}<0.1$ for all $t\in[T]$, since there are at most $10\Delta_{1}$ times when this assumption is violated, contributing $O(\Delta_{1})$ regret. Recall the notation in (16), and we have the following lemma that proves the bids in stage $s$ are only $O(2^{-s})$-suboptimal under $\widehat{r}_{t}$.
@@
In Algorithm 5, for every $t\in[T]$ and $s\in[S]$, it holds that $\max_{b^{*}\in{\mathcal{B}}}\widehat{r}_{t}(b^{*})-\widehat{r}_{t}(b)\leq 8\cdot 2^{-s}$ for all $b\in B_{s}$.
Let $b^{\circ}_{t}$ denote the selected bid before the truncation in Line 22 in Algorithm 5 and $b_{t}$ denote the final bid after. Since Assumption 2.7 implies $\|G_{t}-\widehat{G}_{t}\|_{\infty}\leq 2\delta_{t}$, by Lemma C.4, the truncation step does not hurt the reward too much:
@@
[…]
@@
where (a) is from Lemma 5.5, (b) uses $w^{(s)}_{t}(b_{t}^{\circ})\geq 2^{-s}$ when $t\in\Phi^{(s)}$, and (c) uses $\widehat{G}_{t}(b_{t})(1-\widehat{G}_{t}(b_{t}))\geq\widehat{G}_{t}(b_{t}^{\circ})(1-\widehat{G}_{t}(b_{t}^{\circ}))$ to conclude that $w^{(s)}_{t}(b_{t})\geq w^{(s)}_{t}(b_{t}^{\circ})$ by the definition of $w_{t}^{(s)}$. Then applying the following elliptical potential lemma leads to the upper bound in Theorem 5.4.
@@ 6 Conclusion and Future Directions
Note that Assumption 6.1 does not require the log-concavity assumed in [9], and is satisfied for Gaussian $\eta_{t}$ (by the well-known Mills ratio results [23, Eq (9)]). The following result shows that Assumption 2.7 still holds under the above assumption:
@@
Let Assumptions 2.12.3 and Assumption 6.1 hold. There exists a CDF estimator $\widehat{G}_{t}$ that satisfies Assumption 2.7 with the HOB error bounds $\Delta_{1}={\widetilde{O}}(\sqrt{dT})$ and $\Delta_{2}={\widetilde{O}}(d)$.
Since the quantity $\Delta$ in Assumption 2.5 satisfies $\Delta\leq 2\Delta_{1}$, our algorithms achieve regret ${\widetilde{O}}(\sqrt{dT})$ under Model 2 and ${\widetilde{O}}(d\sqrt{T})$ under Model 3. It is an interesting question to see the tightness of the ${\widetilde{O}}(d\sqrt{T})$ regret under Model 3, and generalize the above results to other models of the HOBs $M_{t}$.
@@ Appendix A Proofs of Main Lemmas
Let $\beta=\sqrt{\log(2T)}$ and $\lambda=1$. Suppose $(v_{s,1},v_{s,0})_{s\in\Phi_{t}}$ are conditionally independent given $(x_{s},b_{s},M_{s})_{s\in\Phi_{t}}$, then with probability at least $1-T^{-2}$, it holds that
@@
We will show the claimed inequality for $i=1$, and the case $i=0$ follows the same argument. Denote $D_{t,1}=[\mathds{1}[b_{\tau}\geq M_{\tau}]x_{\tau}]_{\tau\in\Phi_{t}}\in{\mathbb{R}}^{d\times|\Phi_{t}|}$ and $V_{t,1}=[\mathds{1}[b_{\tau}\geq M_{\tau}]v_{t,1}]_{\tau\in\Phi_{t}}\in{\mathbb{R}}^{1\times|\Phi_{t}|}$. Recall that in LIN-PO we define
@@
Since $(v_{s,1},v_{s,0})_{s\in\Phi_{t}}$ are conditionally independent, we have ${\mathbb{E}}\left[V^{\top}_{t,1}|(x_{s},b_{s},M_{s})_{s\in\Phi_{t}}\right]=D^{\top}_{t,1}\theta_{*,1}$. Thus by Azuma’s inequality [7, 4],
@@
In Algorithm 3, for every $t\in[T]$ and $s\in[S]$, it holds that $\max_{b^{*}\in{\mathcal{B}}}\widehat{r}_{t}(b^{*})-\widehat{r}_{t}(b)\leq 8\cdot 2^{-s}$ for all $b\in B_{s}$.
@@
Suppose the hypothesis holds for $s-1$. The elimination step (Line 3) in Algorithm 3 implies that for any $b\in B_{s}$, we have $w^{(s-1)}_{t}(b)\leq 2^{-(s-1)}$, $w^{(s-1)}_{t}(b^{*}_{t})\leq 2^{-(s-1)}$, and $\widehat{r}_{t}^{(s-1)}(b)+w^{(s-1)}_{t}(b)\geq\widehat{r}_{t}^{(s-1)}(b^{*}_{t})+w^{(s-1)}_{t}(b^{*}_{t})-2^{-s+2}$. Since $b^{*}_{t}\in B_{s-1}$ by the induction hypothesis, by Lemma 4.1,
@@
where (a) uses Lemma 4.1, (b) is by the elimination criterion in Algorithm 5, and (c) follows from $b^{*}_{t}\in B_{s-1}$. This concludes the induction and hence the proof. ∎
A.3 Proof of Lemma 4.5
@@
$\sum_{\tau<t}\|z_{\tau}\|_{A_{\tau}^{-1}}\leq\sqrt{2d(t-1)\log\left(1+\frac{t-1}{d}\right)}.$
@@
[…]
@@
where (a) uses $\det(AB)=\det(A)\det(B)$ and (b) uses that $I+vv^{T}$ only has eigenvalues $1$ and $1+\|v\|_{2}^{2}$. Since $\det(A_{1})=\det(cI)=c$, by AM-GM inequality,
@@
where (a) uses Lemma A.3. Similarly,
@@
A.4 Proof of Lemma 5.1
Lemma A.5 (Restatement of Lemma 5.1).
@@
A.5 Proof of Lemma 5.2
Lemma A.6 (Restatement of Lemma 5.2).
@@
where $\gamma$ is defined in Line 7 of Algorithm 4.
@@
Let the bias of the IPW estimator in (11) at time $\tau$ be $\xi_{\tau}={\mathbb{E}}[\widetilde{e}_{\tau}(b_{\tau})]-\theta_{*}^{\top}x_{\tau}.$ Denote $D_{t}=[\sigma_{\tau}^{-1}x_{\tau}]_{\tau\in\Phi_{t}}\in{\mathbb{R}}^{d\times|\Phi_{t}|}$ the weighted contexts, $V_{t}=[\sigma_{\tau}^{-1}\widetilde{e}_{\tau}(b_{\tau})]_{\tau\in\Phi_{t}}\in{\mathbb{R}}^{|\Phi_{t}|\times 1}$ the weighted estimators, and $\Xi_{t}=[\sigma_{\tau}^{-1}\xi_{\tau}]_{\tau\in\Phi_{t}}\in{\mathbb{R}}^{|\Phi_{t}|\times 1}$ the weighted biases, where $\sigma_{\tau}^{-1}=\sqrt{\widehat{G}_{t}(b_{\tau})\left(1-\widehat{G}_{t}(b_{\tau})\right)}$ as defined in Lemma 5.1. We have
@@
where (b) applies Cauchy-Schwartz inequality and (c) is due to $D_{t}^{\top}(I+D_{t}D_{t}^{\top})D_{t}\preceq I$ by Woodbury matrix identity. To obtain the desired claim, we can further upper bound the following by Lemma 5.1:

### v2

@@ 4.2 Regret Upper Bound
The following theorem establishes the regret upper bound of Algorithm˜3.
Theorem 4.3 (Upper bound of Algorithm˜3).
Let $\pi$ be the resulting bidding strategy of Algorithm˜3 (combined with Algorithm˜2). Then under Assumptions 2.12.3 and 2.5, the policy $\pi$ achieves an expected regret
@@
It is clear that Theorem˜4.3 implies Theorem˜2.6. The remainder of this section is devoted to the proof of Theorem˜4.3. Since all events with probability $O(T^{-1})$ only contribute $O(S)$ regret in total, WLOG we assume that the guarantees in Assumption˜2.5 and Lemma˜4.1 hold almost surely. Adopting the notation [^1] Warning: not to confuse $\widehat{r}_{t}(b)$ with $\widehat{r}_{t}^{(s)}(b)$ in Algorithm 3: $\widehat{r}_{t}(b)$ is the reward w.r.t. $\widehat{G}_{t}(b)$, while $\widehat{r}_{t}^{(s)}(b)$ further plugs in the estimated parameters from stage $s$ for $\theta_{*,0}$ and $\theta_{*,1}$.
@@
In Algorithm˜3, for every $t\in[T]$ and $s\in[S]$, it holds that $\max_{b^{*}\in{\mathcal{B}}}\widehat{r}_{t}(b^{*})-\widehat{r}_{t}(b)\leq 8\cdot 2^{-s}$ for all $b\in B_{s}$.
In other words, Lemma˜4.4 states that the any bid that has survived into stage $s$ is $O(2^{-s})$-close to optimal under $\widehat{r}_{t}$. Consequently, if we denote by $\Phi^{(s)}\subseteq[T]$ the set of time indices belonging to stage $s\in[S]$, and $\Phi^{(S+1)}\subseteq[T]$ the set of time indices where all widths are below $\frac{1}{\sqrt{T}}$, we get
@@
Here (a) uses $|{\mathbb{E}}[r_{t}(b)]-\widehat{r}_{t}(b)|\leq\|G_{t}-\widehat{G}_{t}\|_{\infty}\leq\varepsilon_{t}$ by (2), (10), and Assumption˜2.5, (b) follows from Lemma˜4.4, and (c) uses $S=\lceil\log_{2}\sqrt{T}\rceil$ and $w_{t}^{(s)}(b_{t})\geq 2^{-s}$ whenever $t\in\Phi^{(s)}$. Then Theorem˜4.3 is a direct consequence of the following upper bound, which is essentially the elliptical potential lemma applied to the sum $\sum_{t\in\Phi^{(s)}}w_{t}^{(s)}(b_{t})$.
@@ 5.1 An Adaptive Bid Elimination Algorithm
where $\delta_{t}$ is the concentration parameter in Assumption˜2.7 which is assumed to be known. Unlike $\widehat{e}_{t}(b)$, the truncated estimator $\widetilde{e}_{t}(b)$ is biased. The following lemma summarizes the bias-variance properties of $\widetilde{e}_{t}(b)$.
@@
In the sequel, we will use $\sigma_{t}(b)^{2}$ in (23) as the variance proxy for the truncated IPW estimator $\widetilde{e}_{t}(b)$. In particular, since $\widehat{G}_{t}(b)$ could be close to 0 or $1$ (i.e. no overlap condition), the quantity $\sigma_{t}(b)$ can indeed become prohibitively large.
@@ 5.1.2 Base algorithm
1Input: Time indices $\Phi_{t}\subseteq[t-1]$, bid space $B_{t}\subseteq{\mathcal{B}}$, estimated CDF $\widehat{G}_{t}$, time horizon $T$.
@@
7Set $\gamma\leftarrow\lambda+c_{1}\log(2T)+c_{2}\sqrt{\sum_{\tau\in\Phi_{t}}\delta_{\tau}^{2}}$, with $\delta_{\tau}$ given in Assumption 2.7 and absolute constants $c_{1},c_{2}>0$ defined in Lemma˜A.6.
@@
9   Compute width $w_{t}(b)\leftarrow c_{3}L\gamma\left(\widehat{G}_{t}(b)(1-\widehat{G}_{t}(b))\|x_{t}\|_{A_{t}^{-1}}+\gamma\|x_{t}\|_{A_{t}^{-1}}^{2}\right)$ with absolute constant $c_{3}>0$ defined in Lemma˜A.7.
@@
Next we present the base algorithm LIN-TE in Algorithm˜4 based on $\widetilde{e}_{t}(b)$. Again, we will apply linear regression of $\widetilde{e}_{t}(b_{t})$ against $x_{t}$ to estimate $\theta_{*}$, but Lemma˜5.1 shows that the ordinary least squares will suffer from heteroskedasticity. To address this problem, we use the following weighted least squares estimator:
@@
where $\gamma$ is defined in Line 7 of Algorithm˜4.
As for the quantities $u_{t,0}(b)$ and $u_{t,1}(b)$ in Lines 10 and 11, note that the high-probability event in Lemma˜5.2 yields that for every $b\in{\mathcal{B}}$,
@@ 5.1.3 Master algorithm
1Input: Time horizon $T$, bid space ${\mathcal{B}}=[0,1]$, estimation oracle $\mathcal{O}$}.
2Initialize: set $S=\lceil\log_{2}\sqrt{T}\rceil$, $\Phi^{(s)}_{1}$$\varnothing$ for $s\in[S]$.
@@
Similar to Algorithm˜3, the master algorithm in Algorithm˜5 partitions $[t]$ into multiple stages to ensure the conditional independence condition in Lemma˜5.2. However, the main distinct feature of Algorithm˜5 lies in the selection from two different criteria $u_{t,0}$ and $u_{t,1}$. The reason behind this selection is that, the rewards
@@
in (14) and (15) are sensitive to the estimation error of $\theta_{*}^{\top}x_{t}$ in different ways: in $\widehat{r}_{t,0}(b)$ the leading coefficient is $\widehat{G}_{t}(b)$, while in $\widehat{r}_{t,1}(b)$ the leading coefficient has magnitude $1-\widehat{G}_{t}(b)$. Therefore, a possibly large estimation error of $\theta_{*}^{\top}x_{t}$ when $\widehat{G}_{t}(b)$ is close to 0 (resp. $1$) does not transform into a large estimation error in $\widehat{r}_{t,0}(b)$ (resp. $\widehat{r}_{t,1}(b)$); this is essentially a “variance reduction” step. The key then is to choose the better reward ($\widehat{r}_{t,0}(b)$ or $\widehat{r}_{t,1}(b)$) for decision making, depending on whether $\widehat{G}_{t}(b)$ is close to 0 or $1$ for approximately optimal bids $b$.
To this end, we compute the maximizer $b_{*,1}$ of the map $b\mapsto u_{t,1}(b)$, and check if $\widehat{G}_{t}(b_{*,1})$ is close to 0 or $1$. A small $\widehat{G}_{t}(b_{*,1})$ suggests a small CDF and thus the choice of $\widehat{r}_{t,0}$, and a large $\widehat{G}_{t}(b_{*,1})$ suggests a big CDF and the other option $\widehat{r}_{t,1}$. This is the criterion $i\in\{0,1\}$ in Line 9 of Algorithm˜5. The following result shows that the selected reward $\widehat{r}_{t,i}(b)$ enjoys a small width $w_{t}(b)$ that scales with $\widehat{G}_{t}(b)(1-\widehat{G}_{t}(b))=\sigma_{t}(b)^{-2}$ for all candidate bids.
@@
Suppose the event in Lemma 5.2 holds. Let $\widehat{r}_{t,0}$ and $\widehat{r}_{t,1}$ be defined as in (14) and (15). For any $s\in[S]$ and for the criterion $i\in\{0,1\}$ selected in Algorithm˜5, it holds that
@@
where $b^{(s)}_{*,i}=\operatorname*{arg\,max}_{b\in B_{s}}u^{(s)}_{t,i}(b)$ for $i=0,1$ as defined in Algorithm˜4.
@@ 5.2 Regret Upper Bound
The regret upper bound of Algorithm˜5 is established in the following result, which implies Theorem 2.8.
Theorem 5.4 (Upper bound of Algorithm˜5).
Let $\pi$ be the resulting bidding strategy of Algorithm˜5 (combined with Algorithm˜4). Then under Assumptions 2.12.3 and 2.7, the policy $\pi$ achieves an expected regret
@@
Again, we present the proof of Theorem 5.4 in the remainder of this section. WLOG we assume Assmuption 2.7 and the high-probability event in Lemma˜5.2 holds almost surely. We also assume that $\delta_{t}<0.1$ for all $t\in[T]$, since there are at most $10\Delta_{1}$ times when this assumption is violated, contributing $O(\Delta_{1})$ regret. Recall the notation in (16), and we have the following lemma that proves the bids in stage $s$ are only $O(2^{-s})$-suboptimal under $\widehat{r}_{t}$.
@@
In Algorithm˜5, for every $t\in[T]$ and $s\in[S]$, it holds that $\max_{b^{*}\in{\mathcal{B}}}\widehat{r}_{t}(b^{*})-\widehat{r}_{t}(b)\leq 8\cdot 2^{-s}$ for all $b\in B_{s}$.
Let $b^{\circ}_{t}$ denote the selected bid before the truncation in Line 22 in Algorithm˜5 and $b_{t}$ denote the final bid after. Since Assumption˜2.7 implies $\|G_{t}-\widehat{G}_{t}\|_{\infty}\leq 2\delta_{t}$, by Lemma˜C.4, the truncation step does not hurt the reward too much:
@@
By Assumption 2.7, we can bound both $\text{(a)}\leq 2\sum_{t\in[T]}\delta_{t}=2\Delta_{1}$ and $\text{(e)}\leq 2\Delta_{1}$. By definition of $\widehat{b}^{*}_{t}$, $\text{(b)}\leq 0$. By (17), $\text{(d)}\leq\left(1+c_{1}\log(2T)+c_{2}\sqrt{\Delta_{2}}\right)\sqrt{dT}+8\Delta_{1}$. It remains to bound (c) as follows:
@@
where (a) is from Lemma˜5.5, (b) uses $w^{(s)}_{t}(b_{t}^{\circ})\geq 2^{-s}$ when $t\in\Phi^{(s)}$, and (c) uses $\widehat{G}_{t}(b_{t})(1-\widehat{G}_{t}(b_{t}))\geq\widehat{G}_{t}(b_{t}^{\circ})(1-\widehat{G}_{t}(b_{t}^{\circ}))$ to conclude that $w^{(s)}_{t}(b_{t})\geq w^{(s)}_{t}(b_{t}^{\circ})$ by the definition of $w_{t}^{(s)}$. Then applying the following elliptical potential lemma leads to the upper bound in Theorem˜5.4.
@@ 6 Conclusion and Future Directions
Note that Assumption˜6.1 does not require the log-concavity assumed in [9], and is satisfied for Gaussian $\eta_{t}$ (by the well-known Mills ratio results [23, Eq (9)]). The following result shows that Assumption˜2.7 still holds under the above assumption:
@@
Let Assumptions 2.12.3 and Assumption˜6.1 hold. There exists a CDF estimator $\widehat{G}_{t}$ that satisfies Assumption˜2.7 with the HOB error bounds $\Delta_{1}={\widetilde{O}}(\sqrt{dT})$ and $\Delta_{2}={\widetilde{O}}(d)$.
Since the quantity $\Delta$ in Assumption˜2.5 satisfies $\Delta\leq 2\Delta_{1}$, our algorithms achieve regret ${\widetilde{O}}(\sqrt{dT})$ under Model 2 and ${\widetilde{O}}(d\sqrt{T})$ under Model 3. It is an interesting question to see the tightness of the ${\widetilde{O}}(d\sqrt{T})$ regret under Model 3, and generalize the above results to other models of the HOBs $M_{t}$.
@@
Acknowledgements
This work is generously supported by the NSF award 2106508.
@@ Appendix A Proofs of Main Lemmas
Let $\beta=\sqrt{\log(2T)}$ and $\lambda=1$. Suppose $(v_{\tau,1},v_{\tau,0})_{\tau\in\Phi_{t}}$ are conditionally independent given $(x_{\tau},b_{\tau},M_{\tau})_{\tau\in\Phi_{t}}$, then with probability at least $1-T^{-2}$, it holds that
@@
We will show the claimed inequality for $i=1$, and the case $i=0$ follows the same argument. Denote $D_{t,1}=[\mathds{1}[b_{\tau}\geq M_{\tau}]x_{\tau}]_{\tau\in\Phi_{t}}\in{\mathbb{R}}^{d\times|\Phi_{t}|}$ and $V_{t,1}=[\mathds{1}[b_{\tau}\geq M_{\tau}]v_{\tau,1}]_{\tau\in\Phi_{t}}\in{\mathbb{R}}^{1\times|\Phi_{t}|}$. Recall that in LIN-PO we define
@@
Since $(v_{\tau,1},v_{\tau,0})_{\tau\in\Phi_{t}}$ are conditionally independent, we have ${\mathbb{E}}\left[V^{\top}_{t,1}|(x_{\tau},b_{\tau},M_{\tau})_{\tau\in\Phi_{t}}\right]=D^{\top}_{t,1}\theta_{*,1}$. Thus by Azuma’s inequality [7, 4],
@@
In Algorithm˜3, for every $t\in[T]$ and $s\in[S]$, it holds that $\max_{b^{*}\in{\mathcal{B}}}\widehat{r}_{t}(b^{*})-\widehat{r}_{t}(b)\leq 8\cdot 2^{-s}$ for all $b\in B_{s}$.
@@
Suppose the hypothesis holds for s1s-1. The elimination step (Line 3) in Algorithm˜3 implies that for any $b\in B_{s}$, we have $w^{(s-1)}_{t}(b)\leq 2^{-(s-1)}$, $w^{(s-1)}_{t}(b^{*}_{t})\leq 2^{-(s-1)}$, and $\widehat{r}_{t}^{(s-1)}(b)+w^{(s-1)}_{t}(b)\geq\widehat{r}_{t}^{(s-1)}(b^{*}_{t})+w^{(s-1)}_{t}(b^{*}_{t})-2^{-s+2}$. Since $b^{*}_{t}\in B_{s-1}$ by the induction hypothesis, by Lemma˜4.1,
@@
where (a) uses Lemma˜4.1, (b) is by the elimination criterion in Algorithm˜5, and (c) follows from $b^{*}_{t}\in B_{s-1}$. This concludes the induction and hence the proof. ∎
A.3 Proof of Lemma˜4.5
@@
$\sum_{\tau<t}\|z_{\tau}\|_{A_{\tau}^{-1}}\leq\sqrt{2d(t-1)\log\left(1+\frac{t-1}{d}\right)}+\sqrt{2(t-1)\log(c)}.$
@@
$\overset{\text{(a)}}{=}\det\left(A_{t-1}\right)\det\left(I+A_{t-1}^{-1/2}z_{t-1}z_{t-1}^{\top}A_{t-1}^{-1/2}\right)$
$\overset{\text{(b)}}{=}\det\left(A_{\tau}\right)\left(1+\|z_{t-1}\|_{A_{t-1}^{-1}}^{2}\right)$
@@
where (a) uses $\det(AB)=\det(A)\det(B)$ and (b) uses that $I+vv^{T}$ only has eigenvalues 11 and $1+\|v\|_{2}^{2}$. Since $\det(A_{1})=\det(cI)=c$, by AM-GM inequality,
@@
where (a) uses Lemma˜A.3. Similarly,
@@
A.4 Proof of Lemma˜5.1
Lemma A.5 (Restatement of Lemma˜5.1).
@@
A.5 Proof of Lemma˜5.2
Lemma A.6 (Restatement of Lemma˜5.2).
@@
where $\gamma$ is defined in Line 7 of Algorithm˜4.
@@
Let the bias of the IPW estimator in (11) at time $\tau$ be $\xi_{\tau}={\mathbb{E}}[\widetilde{e}_{\tau}(b_{\tau})]-\theta_{*}^{\top}x_{\tau}.$ Denote $D_{t}=[\sigma_{\tau}^{-1}x_{\tau}]_{\tau\in\Phi_{t}}\in{\mathbb{R}}^{d\times|\Phi_{t}|}$ the weighted contexts, $V_{t}=[\sigma_{\tau}^{-1}\widetilde{e}_{\tau}(b_{\tau})]_{\tau\in\Phi_{t}}\in{\mathbb{R}}^{|\Phi_{t}|\times 1}$ the weighted estimators, and $\Xi_{t}=[\sigma_{\tau}^{-1}\xi_{\tau}]_{\tau\in\Phi_{t}}\in{\mathbb{R}}^{|\Phi_{t}|\times 1}$ the weighted biases, where $\sigma_{\tau}^{-1}=\sqrt{\widehat{G}_{t}(b_{\tau})\left(1-\widehat{G}_{t}(b_{\tau})\right)}$ as defined in Lemma˜5.1. We have
@@
where (b) applies Cauchy-Schwartz inequality and (c) is due to $D_{t}^{\top}(I+D_{t}D_{t}^{\top})^{-1}D_{t}\preceq I$ by Woodbury matrix identity. To obtain the desired claim, we can further upper bound the following by Lemma 5.1: